HETAV PATEL

Grade 11,

A car travelling at 60 km/h overtakes another car travelling at 42 km/h. Assuming each car to be 5.0 m long, find the time taken during the overtake and the total road distance used for the overtake.

A car travelling at 60 km/h overtakes another car travelling at 42 km/h. Assuming each car to be 5.0 m long, find the time taken during the overtake and the total road distance used for the overtake.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. v1 = 60 km/hr = 16.6 m/s. v2 = 42 km/h = 11.6 m/s. Relative velocity between the cars = (16.6 – 11.6) = 5 m/s. Distance to be travelled by first car is 5 + t = 10 m. Time = t = s/v = 0/5 = 2 sec to cross the 2nd car. In 2 sec the 1st car moved = 16.6 × 2 = 33.2 m H also covered its own length 5 m. ∴ Total road distance used for the overtake = 33.2 + 5 = 38 m.